# Spectral properties from an efficient analytical representation of the $GW$ self-energy within a multipole approximation

### v1

@@ III.1 Self-energy and Green’s function of prototypical materials
Panels (g-i) of Fig. 4 show the renormalization factor corresponding to the reference FF data evaluated with Eq. (11) (red squares) and the MPA-$\Sigma$ representation of Eq. (13) (blue circles). In both cases, $Z$ is calculated from the local derivative of $\Sigma$, according to Eq. (7). They differ by less than 0.003, 0.010 and 0.025 for all the quasiparticles of Si, Na, and Cu, respectively. The results labeled MPA-$G$ (yellow triangles) were obtained with the definition proposed in Eq. (16), in which $Z$ is given by the largest residue of $G$. For most of the quasiparticles, $Z^{\text{MPA-}\Sigma}$ is quite similar to $Z^{\text{MPA-}G}$; however, for the Na and Cu states with more intense satellites ($\mathrm{Re}~{}Z\lesssim 0.8$), the deviation can increase to 0.045 and 0.070 respectively, which states that these two $Z$ definitions differ on a fundamental level. A detailed analysis of their differences is presented in Sec. III.3.
@@
Fig. 5 shows the QP band structure of Na (a), Si (b), monolayer MoS2 (c), and Cu (d). The width of the lines gives the imaginary part of the QP poles $\mathrm{Im}~{}[\epsilon_{n{\mathbf{k}}}]$, while the color shade indicates the value of the renormalizaion factors $\mathrm{Re}~{}[Z_{n{\mathbf{k}}}]$ for valence (orange shades) and conduction (purple shades) bands. In general, we find that both $\mathrm{Im}~{}[\epsilon_{n{\mathbf{k}}}]$ and $\mathrm{Re}~{}[Z_{n{\mathbf{k}}}]$ increase further away from the Fermi level, as a result of the QP pole and the satellites broadening and merging in a single peak. However, in the energy range in which multiple bands of different character cross, the picture becomes more complex showing a non-monotonic behavior. The metallic band of Na is an exception, with $\mathrm{Im}~{}[\epsilon_{n{\mathbf{k}}}]$ and $\mathrm{Re}~{}[Z_{n{\mathbf{k}}}]$ decreasing without crossing any other band. This is in line with the increased weight of the satellites, as discussed in Sec. III.2.
Figure 6: Spectral self-energy (panels (a-d)) and Green’s function (panels (e-h)) band structures of Na, for the first valence state (a, e), the metallic band (b, f), the first two valence bands (c, g) and their combinations (d, h).
Figure 7: Spectral self-energy (a-c) and Green’s function (d-f) band structure of Si (left panels), MoS2 (central panels), Cu (right panels). We considered the first 4 valence and 6 conduction bands of Si along the $L\Gamma X$ path, 9 valence and 7 conduction bands of MoS2 along the $\Gamma MK\Gamma$ path, and 6 valence and 9 conduction bands of Cu along the $L\Gamma X$ path.
@@ III.2 Spectral band structures
The spectral function probed in photoemission and inverse photoemission experiments is the sum of the spectral function of each state, corresponding respectively to valence and conduction bands, according to the polarization of the incoming light [8]. In order to simultaneously represent an uneven number of valence and conduction states, we have defined the following spectral functions of the self-energy and the Green’s function:
$\mathrm{Im}~{}\Sigma_{{\mathbf{k}}}(\omega)\equiv\frac{1}{n_{v}}\sum_{v}^{n_{v}}\mathrm{Im}~{}\Sigma_{v{\mathbf{k}}}(\omega)+\frac{1}{n_{c}}\sum_{c}^{n_{c}}\mathrm{Im}~{}\Sigma_{c{\mathbf{k}}}(\omega)$ (17)
$\mathrm{Im}~{}G_{{\mathbf{k}}}(\omega)\equiv\frac{1}{n_{v}}\sum_{v}^{n_{v}}\mathrm{Im}~{}G_{v{\mathbf{k}}}(\omega)+\frac{1}{n_{c}}\sum_{c}^{n_{c}}\mathrm{Im}~{}G_{c{\mathbf{k}}}(\omega),$ (18)
where $v$ and $c$ run over valence and conduction states respectively and $n_{v}$ and $n_{c}$ correspond to the total number of valence and conduction bands considered. Since we use time-ordered operators, the spectra have opposite signs for valence and conduction states. We include a normalization based on the number of valence and conduction bands to balance the background intensity of the two regions, and facilitate its visualization. In a detailed comparison with experiment, it would be necessary to properly account for the intensity resulting from the dipole selection rules dictated by the light polarization [8].
Fig. 6 shows $\mathrm{Im}~{}\Sigma_{{\mathbf{k}}}$ (a-d) and $\mathrm{Im}~{}G_{{\mathbf{k}}}$ (e-h) of Na along the $\Gamma N$ ${\mathbf{k}}$-path, for the highest non-metallic valence band (v1), for the metallic band that crosses the Fermi level close to $N$ (m1), the two lowest conduction bands (c1-c2), and the combination of all these 4 bands (all). The intensity of the spectra ranges from positive values (orange shades), corresponding to a valence band character, to negative (purple shades), indicating a conduction band character.
@@
$\mathrm{Im}~{}\Sigma_{{\mathbf{k}}}$ exhibits bands that arise from the coupling of the plasmon with the single KS states, as discussed in the previous section. From now on we will call them $\Sigma$ bands. Essentially, the position of the $\Sigma$ valence (conduction) bands corresponds to the independent-particle bands of $G_{0}$ shifted down (up) by the plasmon energy, $\zeta_{n{\mathbf{k}}}\sim\varepsilon^{\text{KS}}_{n{\mathbf{k}}}\pm\omega_{\text{pl}}$, where $\omega_{\text{pl}}=5.8$ eV [40]. Due to the time ordering, each valence and conduction state has a $\mathrm{Im}~{}[\Sigma_{n{\mathbf{k}}}](\omega)$ component with the opposite sign discernible, e.g., in panel (b). Therefore, the spectral contribution of each state to the total spectral function $\mathrm{Im}~{}[\Sigma_{{\mathbf{k}}}]$ is not always trivial, as in the case of the conduction bands in Fig. 6 (d).
@@
$\mathrm{Im}~{}G_{{\mathbf{k}}}$ (panels (e-h)) also exhibits bands that we will call $G$ bands. There are two types, those coming from the QP pole, $\varepsilon^{\text{QP}}_{n{\mathbf{k}}}$, and those formed from the satellites, $\varepsilon^{\text{sat}}_{n{\mathbf{k}}}$. The QP bands are shifted with respect to the $G_{0}$ bands by the $G_{0}W_{0}$ correction, according to Eq. (5). The satellite bands (called sidebands in Ref. [4]) correspond to replicas of the QP bands located at larger energies. As already seen in Fig. 2 (g-i), the QP peak typically dominates the spectral function. To better discern the satellite structures, we impose a threshold of $|\mathrm{Im}~{}G_{{\mathbf{k}}}|\leq 0.8$ eV-1 in the color map. In the case of the metallic band of Na, the intensity of the satellite around the $\Gamma$ point is similar in magnitude to the QP peak, which is interpreted as a plasmaron [75, 4, 76, 5].
Since the satellites emerge from the plasmonic structures in $\Sigma$ [5], the satellite bands are shifted with respect to the QP bands by roughly the energy of the plasmon [19], $\varepsilon^{\text{sat}}_{n{\mathbf{k}}}\sim\varepsilon^{\text{QP}}_{n{\mathbf{k}}}\pm\omega_{\text{pl}}$, and, in turn, are shifted with respect to the $\Sigma$ bands by the $G_{0}W_{0}$ correction, $\varepsilon^{\text{sat}}_{n{\mathbf{k}}}-\zeta_{n{\mathbf{k}}}\sim\varepsilon^{\text{QP}}_{n{\mathbf{k}}}-\varepsilon^{\text{KS}}_{n{\mathbf{k}}}$. The connection between the poles of $\Sigma$ and $G$ is illustrated with the toy model introduced in Sec. III.3. Due to this connection, the accuracy of both the $\Sigma$ and $G$ bands requires a good description of the screening, which can be improved with vertex corrections [75, 19] or cumulant expansions [19, 76, 20, 14]. At finite temperature, electron-phonon interactions are expected to further renormalize and broaden the QP peak [21, 14], as shown for example in Refs. [77, 10, 78] for the case of Cu.
Fig. 7 shows $\mathrm{Im}~{}\Sigma_{{\mathbf{k}}}$ (top) and $\mathrm{Im}~{}G_{{\mathbf{k}}}$ (bottom) of the first 4 valence and 6 conduction bands of Si along the $L\Gamma X$ path (a, d), 9 valence and 7 conduction bands of MoS2 along the $\Gamma MK\Gamma$ path (b, d), and 6 valence and 9 conduction bands of Cu along the $L\Gamma X$ path (c, f). The color maps for the intensity of ${\mathrm{Im}~{}}G_{{\mathbf{k}}}$ have thresholds of ±$\pm 0.3$, ±$\pm 0.5$, and ±$\pm 0.7$ eV-1 for Si, MoS2, and Cu respectively.
Similar to the case of Na, the spectral functions of Si, $\mathrm{Im}~{}\Sigma_{{\mathbf{k}}}$ (a) and $\mathrm{Im}~{}G_{{\mathbf{k}}}$ (d), exhibit well defined bands, with $\Sigma$ and satellite bands respectively shifted from the $G_{0}$ and QP bands by the plasmon energy, $\omega_{\text{pl}}=16.6$ eV [39, 19]. This results in a $\Sigma$ band gap given by the KS gap plus twice the plasmon energy ($\varepsilon_{\text{gap}}^{\Sigma}=\varepsilon_{\text{gap}}^{\text{KS}}+2\omega_{\text{pl}}$).
@@
The $\Sigma$ bands of MoS2 and Cu (Fig. 7 (b) and (c)) are generally broader and overlap more with each other than those of Na and Si. This is in line with both the more complex structure of $\Sigma$ and the fact that the energy separation of the bands is smaller than the width of the main plasmon, located at 10.98 and 26.5 eV for MoS2 and Cu respectively. As a result, MoS2 exhibits rather flat and broadened valence and conduction $\Sigma$ bands, with an apparent gap given by secondary peaks at energies smaller than the main plasmon. The valence of Cu also shows a main flat and broadened dispersion, even if some secondary bands are still visible. As a consequence, the satellites of MoS2 and Cu are also broadened, resulting in the diffused background of Fig. 7 (e, f).
III.3 The quasiparticle picture in MPA models
@@
Figure 8: Real and imaginary parts of the scaled poles (a, c) and residues (b, d) of the Green’s function model of Eq. (21) as a function of the $Z$ parameter, where we have fixed $x=0$. The solid horizontal gray line in panel (a) corresponds to $\mathrm{Re}~{}[\epsilon_{p}/\xi]=1$, while a dashed line is drawn at zero. In panel (b), the gray dashed horizontal line at $\mathrm{Re}~{}[P_{p}]=0.5$ represents the limit of the QP picture, while the solid line at 1 indicates the sum rule of Eq. (15) ($P_{1}+P_{2}=1$). The dotted black line corresponds to $\mathrm{Re}~{}[Z^{\text{lin}}]=\mathrm{Re}~{}Z$, and by contrasting it with $\mathrm{Re}~{}[Z^{\text{MPA-}G}_{\text{s1}}]=\mathrm{Re}~{}[P_{1}]$ we are comparing the two definitions of the renormalization factor of Eq. (7) and Eq. (16).
As shown in the previous sections, the self-energy can be approximated by an MPA-$\Sigma$ representation with a small number of poles, from which an MPA-$G$ representation is obtained. Here, we introduce a toy MPA-$\Sigma$ model as a means to analyze the QP picture. We compare the renormalization factor computed from the local derivative of $\Sigma$, as defined in Eq. (7), with the definition in an MPA-$G$ representation proposed in Eq. (16).
@@
$\Sigma^{\text{s1}}(\omega)=x\xi+\frac{1-Z}{Z}\frac{\xi^{2}}{\omega-\xi},$ (19)
where $\omega$ is centered on the given QP energy, $x\xi$ accounts for static contributions such as the exchange interaction and vertex corrections, and the second term accounts for correlation with a single pole at the plasmon energy $\xi$. The residue of the pole, $S=\xi^{2}(1-Z)/Z$, is defined in terms of a parameter $Z$ in order to coincide with the local definition of the renormalization factor in Eq. (7), independently of the other two parameters:
$Z^{\text{lin}}_{\text{s1}}\equiv\left[1-\frac{\partial\Sigma^{\text{s1}}(\omega)}{\partial\omega}\bigg{|}_{\omega=0}\right]^{-1}=Z.$ (20)
@@
$G^{\text{s1}}(\omega)=\sum_{p=1}^{2}\frac{P^{\text{s1}}_{p}(x,Z)}{\omega-\epsilon^{\text{s1}}_{p}(x,Z,\xi)},$ (21)
which has 2 poles, $\epsilon^{\text{s1}}_{p}$, with residues $P^{\text{s1}}_{p}$, given by
$\epsilon^{\text{s1}}_{1,2}&=\frac{\xi}{2}\left(1+x\pm\sqrt{D}\right)\\ P^{\text{s1}}_{1,2}&=\frac{1}{2}\left(1\pm\frac{1-x}{\sqrt{D}}\right),\end{aligned}$ (22)
@@
$D\equiv(1-x)^{2}+4\frac{1-Z}{Z}.$ (23)
The poles are proportional to $\xi$ and can be rescaled to obtain dimensionless units. The renormalization factor in the MPA-$G$ representation analogous to Eq. (16), corresponds to the largest residue of $G^{\text{s1}}$:
$Z^{\text{MPA-}G}_{\text{s1}}=\left\{\begin{aligned} &P^{\text{s1}}_{1},\qquad\mathrm{Re}~{}[P^{\text{s1}}_{1}]\geq\mathrm{Re}~{}[P^{\text{s1}}_{2}]\\ &P^{\text{s1}}_{2},\qquad\mathrm{Re}~{}[P^{\text{s1}}_{1}]<\mathrm{Re}~{}[P^{\text{s1}}_{2}]\end{aligned}\right.$ (24)
@@
Figure 9: Real and imaginary parts of the scaled poles (a, c) and residues (b, d) of the Green’s function model of Eq. (21) as a function of the $Z$ parameter, where we have fixed its imaginary part to a small value of $\mathrm{Im}~{}Z=0.01$. The solid and dashed horizontal gray lines and the black dotted line in panel (b) are analogous to Fig. 8.
@@
We first focus on the correlation effects by setting $x=0$. Notice that $Z=1$ corresponds to $S=0$ (zero correlation), while $Z\to 0$ corresponds to the limit of infinite correlation. Fig. 8 shows the real and imaginary parts of the scaled poles, $\epsilon^{\text{s1}}_{p}/\xi$ (panels a, c), and the residues, $P_{p}^{\text{s1}}$ (panels b, d), of $G^{\text{s1}}$ as functions of $\mathrm{Re}~{}Z$, for $\mathrm{Im}~{}Z=0$ (blue curves) and $\mathrm{Im}~{}Z=0.3$ (orange curves). Since $x=0$, $\mathrm{Re}~{}Z=1$ and $\mathrm{Im}~{}Z=0$ also correspond to the independent-particle limit, in which the QP state has a zero self-energy correction ($\epsilon^{\text{s1}}_{1}=0$) and carries the whole spectral weight ($P^{\text{s1}}_{1}=Z=1$). As a consequence, the satellite $\epsilon^{\text{s1}}_{2}=\xi$ vanishes ($P^{\text{s1}}_{2}=0$).

### v2

@@ III.1 Self-energy and Green’s function of prototypical materials
Figures 4(g)-4(i) show the linearized renormalization factor of Eq. (7), $Z^{\text{lin}}$, corresponding to the reference FF data evaluated with Eq. (12) (red squares) and the MPA-$\Sigma$ representation of Eq. (14) (blue circles). They differ by less than 0.003, 0.010, and 0.025 for all the quasiparticles of Si, Na, and Cu, respectively. The results labeled MPA-$G$ (yellow triangles) correspond to the non-linearized $Z$, obtained as the residue $Z_{p}$ of the QP pole in Eqs. (22) and (23). Therefore, comparing the MPA-$\Sigma$ and MPA-$G$ results corresponds to comparing $Z^{\text{lin}}$ with $Z$. For most of the quasiparticles, the linearization is a good approximation and $Z^{\text{MPA-}\Sigma}$ is quite similar to $Z^{\text{MPA-}G}$; however, for the Na and Cu states with more intense satellites ($\mathrm{Re}~{}Z\lesssim 0.8$), the deviation can increase to 0.045 and 0.070, respectively. A detailed analysis of the limitations of Eq. (7) as an approximation of the spectral weight of the QP pole in different scenarios is presented in Appendix A. In particular, aside from the QP correction to the single-particle energies, it is shown that the deviation can increase with the QP broadening.
@@
Figure 5 shows the QP band structure of Na (a), Si (b), monolayer MoS2 (c), and Cu (d), interpolated in ${\mathbf{k}}$-space, as described in Appendix B. The width of the lines gives the imaginary part of the QP poles $\mathrm{Im}[\epsilon_{n{\mathbf{k}}}]$, while the color shade indicates the value of the renormalization factors $\mathrm{Re}[Z_{n{\mathbf{k}}}]$ for valence (orange shades) and conduction (purple shades) bands. In general, we find that both $\mathrm{Im}[\epsilon_{n{\mathbf{k}}}]$ and $\mathrm{Re}[Z_{n{\mathbf{k}}}]$ increase further away from the Fermi level, as a result of the QP pole and the satellites broadening and merging in a single peak. However, in the energy range in which multiple bands of different character cross, the picture becomes more complex showing a non-monotonic behavior. The metallic band of Na is an exception, with $\mathrm{Im}[\epsilon_{n{\mathbf{k}}}]$ and $\mathrm{Re}[Z_{n{\mathbf{k}}}]$ decreasing without crossing any other band. This is in line with the increased weight of the satellites, as discussed in Sec. III.2.
Figure 6: Spectral band structures $A_{\Sigma}({\mathbf{k}},\omega)$ (a)-(d) and $A_{G}({\mathbf{k}},\omega)$ (e)-(h) of Na (a), (e); Si (b), (f); MoS2 (c), (g); and Cu (d), (h); computed with MPA-$\Sigma$ and MPA-$G$. We considered the metallic band and the next valence and 2 conduction bands of Na along the $\Gamma N$ path, 4 valence and 6 conduction bands of Si along the $L\Gamma X$ path, 9 valence and 7 conduction bands of MoS2 along the $\Gamma MK\Gamma$ path, and 6 valence and 9 conduction bands of Cu along the $L\Gamma X$ path.
@@ III.2 Spectral band structures
The spectral functions probed in photoemission and inverse photoemission experiments account for the contributions of each state, according to the polarization of the incoming light [8]. In a typical ab initio calculation, the total $\Sigma$ and $G$ spectral functions, $A_{\Sigma}({\mathbf{k}},\omega)\equiv 1/\pi\sum_{n}\mathrm{Im}[\Sigma_{n{\mathbf{k}}}(\omega)]$ and $A_{G}({\mathbf{k}},\omega)\equiv 1/\pi\sum_{n}\mathrm{Im}[G_{n{\mathbf{k}}}(\omega)]$, are evaluated with a finite number of bands. Since we use time-ordered operators, the spectra have opposite signs for valence and conduction states, which makes the background intensity sensitive to the number of bands included. In Sec. III of Ref. [73] we provide detailed descriptions on how we plot spectral functions, especially when few bands are included.
Fig. 6 shows the computed $A_{\Sigma}({\mathbf{k}},\omega)$ (top panels) and $A_{G}({\mathbf{k}},\omega)$ (bottom panels) spectral band structures of Na [Figs. 6(a) and 6(e)], Si [Figs. 6(b) and 6(f)], monolayer MoS2 [Figs. 6(c) and 6(g)] and Cu [Figs. 6(d) and 6(h)], obtained with MPA-$\Sigma$ and MPA-$G$. Their accuracy is comparable with the spectral functions obtained with a FF evaluation (see Sec. III of Ref. [73]). As in Fig. 5, to build such spectral band structures we use the spline interpolation in ${\mathbf{k}}$-space detailed in Appendix B. Due to the time ordering, each state has positive (orange shades) and negative (purple shades) $1/\pi\mathrm{Im}[\Sigma_{n{\mathbf{k}}}(\omega)]$ components, even if positive (negative) intensities are predominant for valence (conduction) states. Therefore, the spectral contribution of each state to the total spectral function $A_{\Sigma}({\mathbf{k}},\omega)$ is not always trivial, as in the case of Na (see Fig. S2 of Ref. [73]).
@@
$A_{\Sigma}({\mathbf{k}},\omega)$ exhibits bands that arise from the coupling of the plasmon with the single KS states, as discussed in the previous section. From now on we will call them $\Sigma$ bands. Essentially, the position of the $\Sigma$ valence (conduction) bands corresponds to the independent-particle bands of $G_{0}$ shifted down (up) by the plasmon energy, $\zeta_{n{\mathbf{k}}}\sim\varepsilon^{\text{KS}}_{n{\mathbf{k}}}\pm\omega_{\text{pl}}$, where $\omega_{\text{pl}}$ has a value of $5.8$, $16.6$, $10.98$, and $26.5$ eV for Na, Si, MoS2, and Cu respectively. Therefore, in semiconducting materials like Si, the $\Sigma$ bands present a gap given by the KS gap plus twice the plasmon energy ($\varepsilon_{\text{gap}}^{\Sigma}=\varepsilon_{\text{gap}}^{\text{KS}}+2\omega_{\text{pl}}$).
@@
$A_{G}({\mathbf{k}},\omega)$ (panels (e-h)) also exhibits bands that we will call $G$ bands. There are two types, those coming from the QP pole, $\varepsilon^{\text{QP}}_{n{\mathbf{k}}}$, and those formed from the satellites, $\varepsilon^{\text{sat}}_{n{\mathbf{k}}}$. The QP bands are shifted with respect to the $G_{0}$ bands by the $G_{0}W_{0}$ correction, according to Eq. (5). The satellite bands (called sidebands in Ref. [4]) correspond to replicas of the QP bands located at larger energies. As already seen in Figs. 2(g)-2(i), the QP peak typically dominates the spectral function. To better discern the satellite structures, we impose thresholds to the $A_{G}({\mathbf{k}},\omega)$ color maps, corresponding to ±$\pm 0.25$, ±$\pm 0.10$, ±$\pm 0.16$, and ±$\pm 0.22$ eV-1 for Na, Si, MoS2, and Cu, respectively. In the case of the metallic band of Na, the intensity of the satellite around the $\Gamma$ point is similar in magnitude to the QP peak, which is interpreted as a plasmaron [81, 4, 82, 5]. Since the satellites emerge from the plasmonic structures in $\Sigma$ [5], the satellite bands are shifted with respect to the QP bands by roughly the energy of the plasmon [21], $\varepsilon^{\text{sat}}_{n{\mathbf{k}}}\sim\varepsilon^{\text{QP}}_{n{\mathbf{k}}}\pm\omega_{\text{pl}}$, and, in turn, are shifted with respect to the $\Sigma$ bands by the $G_{0}W_{0}$ correction, $\varepsilon^{\text{sat}}_{n{\mathbf{k}}}-\zeta_{n{\mathbf{k}}}\sim\varepsilon^{\text{QP}}_{n{\mathbf{k}}}-\varepsilon^{\text{KS}}_{n{\mathbf{k}}}$. One of the advantages of the MPA-$G$ representation is the possibility to analytically separate the spectral contributions of the QP pole and the satellites, as done in Fig. S4 of Ref. [73].
@@
Figure 6 exhibits a qualitative difference between the $\Sigma$ bands of Na [Fig. 6(a)] and Si [Fig. 6(b)], which are easy to isolate, compared to those of MoS2 [Fig. 6(c)] and Cu [Fig. 6(d)], which are generally broader and overlap more with each other. This is in line with both the more complex $\Sigma$ structure of MoS2 and Cu, and the fact that the energy separation of the bands is smaller than the width of the main plasmon. As a result, MoS2 exhibits rather flat and broadened valence and conduction $\Sigma$ bands, with an apparent gap given by secondary peaks at energies smaller than the main plasmon. The $\Sigma$ bands of Cu also shows a main flat and broadened dispersion, even if some secondary bands are still visible. As a consequence, the satellites bands of MoS2 and Cu are also broadened, resulting in the diffuse background of Fig. 6 (g, h).
The results in Fig. 6 illustrate the connection between the poles of $\Sigma$ and $G$, whose analyses allow to identify the QP pole from the satellites, as also illustrated with the toy models introduced in Appendix A. Due to this connection, the accuracy of both the $\Sigma$ and $G$ bands requires a good description of the screening, which can be improved with vertex corrections [81, 21] or cumulant expansions [21, 82, 22, 16]. At finite temperature, electron-phonon interactions are expected to further renormalize and broaden the QP peak [23, 16], as shown for example in Refs. [83, 12, 84] for the case of Cu.
IV Conclusions
We have presented MPA-$\Sigma$, a robust method to efficiently approximate the frequency dependence of the $GW$ self-energy as a multipole-Padé representation, typically with around 10 poles. Such representation, similar to the multipole approximation for the screening interaction, MPA-$W$, is built from numerical data evaluated on around 20 frequency points in the complex plane, thus avoiding explicit evaluations of the self-energy on dense frequency grids (of the order of 1000 frequencies in our cases). MPA-$\Sigma$ allows also to solve the QP equation analytically and obtain an MPA-$G$ representation of the interacting Green’s function, from which all the spectral properties can be easily extracted, including the positions, broadenings, and the spectral weights of the QP pole and its satellites.
Combining MPA-$W$ and MPA-$\Sigma$ is a computationally powerful approach, reducing the number of frequency evaluations by a factor $\sim 50^{2}$, compared to full-frequency approaches, while providing spectra with comparable numerical accuracy. The excellent accuracy of this method has been verified for several materials: bulk Si, Na, and Cu, monolayer MoS2, the NaCl ion pair, and the F2 molecule. The efficiency of the MPA method allows us to compute $\Sigma$ and $G$ spectra in a wide energy range. In particular, our results for NaCl and F2 exhibit features beyond the typical energy range of the $GW$ spectra found in the literature for such molecular species. We have also presented a method for interpolating the spectra of higher-dimensional systems in momentum space, and construct $\Sigma$ and $G$ spectral band structures. For Na, Si, monolayer MoS2, and Cu, we report full $\Sigma$ and $G$ spectral band structures, while isolating the contributions from the QP pole and the satellites.
The spectral weights of both the QP pole and the satellites computed with MPA-$G$ comply with the sum rule for the number of particles and holes, providing an accurate way to evaluate the renormalization factor beyond the linearized QP equation. Therefore, such analytical representations can be useful in understanding the physical nature of the QP picture and the renormalization factor. In the following Appendixes, we present toy models that capture the most typical situations when solving the QP equation in prototypical materials, exposing the limitations of its linearization in different regimes from weak to strong correlation.
Acknowledgments
This work was funded by the Research Council of Norway through the MORTY project (315330). Access to high performance computing resources was provided by UNINETT Sigma2 (NN9711K) in Norway, and by EuroHPC Joint Undertaking through the project EHPC-EXT-2022E01-022 that grants access to Leonardo-Booster@Cineca, Italy. We acknowledge Andrea Ferretti for insightful discussions and comments on the manuscript, and Kristian S. Thygesen and Mikael Kuisma for stimulating discussions.
V Data Availability Statements
The data generated in this article are openly available [85].
Appendix A The quasiparticle picture in the MPA representation
As discussed in Sec. II, it is well known that the linearized renormalization factor in Eq. (7) is not always a good approximation for the QP spectral weight, while there is no established method to compute the spectral weight of satellites. Equation (7) is expected to work when the QP correction is small, although it is still widely used in ab initio calculations even in the case of large corrections, especially in high-throughput studies (see, e.g., Ref. [67]). In the following sections, we expose in detail some of the limitations of such linearization, while highlighting the advantages of the MPA-$G$ representation, consistent with Eqs. (23) and (24), to analyze the QP picture in the limit of weak and strong correlation. To that aim, we use toy MPA-$\Sigma$ models in which the number of poles is limited to one, $\Sigma^{\text{s1}}$, and two, $\Sigma^{\text{s2}}$, and their corresponding MPA-$G$ solutions, $G^{\text{s1}}$ and $G^{\text{s2}}$. Such models, although simple, still capture the most typical situations when solving the QP equation in prototypical materials.
In the next sections we show that (1) the linearized QP equation can incorrectly predict values of $Z^{\text{lin}}<0.5$ even when the QP picture holds ($Z>0.5$). (2) The error introduced by the linearization, not only increases with the magnitude of the QP correction, but also with the QP broadening. (3) The static term in the self-energy plays an important role in the QP picture, affecting the distribution of the spectral weight between the QP pole and the satellites. For certain conditions, the satellites can have a larger spectral weight than the QP pole. (4) In the case of a self-energy with multiple poles, the linearized $Z^{\text{lin}}$ can deviate from the exact value $Z$, even in the limit $Z^{\text{lin}}\to 1$, which does not necessarily correspond to a vanishing linearized QP correction.
A.1 Toy MPA-$\Sigma$ model with a single pole
@@
Figure 7: Real and imaginary parts of the scaled poles (a), (c) and residues (b), (d) of the Green’s function model of Eq. (27) as a function of the $\lambda$ parameter, where we have fixed $x=0$. The solid horizontal gray line in (a) corresponds to $\mathrm{Re}[\epsilon_{p}/\xi]=1$, while a dashed line is drawn at zero. In (b), the gray dashed horizontal line at $\mathrm{Re}[Z_{p}]=0.5$ represents the limit of the QP picture, while the gray solid line at 1 indicates the sum rule of Eq. (24) ($Z_{1}+Z_{2}=1$). The dotted black line corresponds to $\mathrm{Re}[Z^{\text{lin}}]=1-\mathrm{Re}~{}\lambda$, contrasted with $\mathrm{Re}[Z^{\text{MPA-}G}_{\text{s1}}]=\mathrm{Re}[Z_{1}]$.
@@
$\Sigma^{\text{s1}}(\omega)=x\xi+\frac{\lambda}{1-\lambda}\frac{\xi^{2}}{\omega-\xi},$ (25)
where $\omega$ is centered on the given QP energy, $x\xi$ accounts for static contributions such as the exchange interaction and vertex corrections, and the second term accounts for correlation with a single pole at the plasmon energy $\xi$. The residue of the pole, $S=\xi^{2}\lambda/(1-\lambda)$, is defined in terms of a parameter $\lambda$ so that the linearized renormalization factor in Eq. (7), independently of the other two parameters, is given by:
$Z^{\text{lin}}_{\text{s1}}\equiv\left[1-\frac{\partial\Sigma^{\text{s1}}(\omega)}{\partial\omega}\bigg{|}_{\omega=0}\right]^{-1}=1-\lambda.$ (26)
@@
$G^{\text{s1}}(\omega)=\sum_{p=1}^{2}\frac{Z^{\text{s1}}_{p}(x,\lambda)}{\omega-\epsilon^{\text{s1}}_{p}(x,\lambda,\xi)},$ (27)
which has two poles, $\epsilon^{\text{s1}}_{p}$, with residues $Z^{\text{s1}}_{p}$, given by
$\epsilon^{\text{s1}}_{1,2}&=\frac{\xi}{2}\left(1+x\mp\sqrt{D}\right),\\ Z^{\text{s1}}_{1,2}&=\frac{1}{2}\left(1\pm\frac{1-x}{\sqrt{D}}\right),\end{aligned}$ (28)
@@
$D\equiv(1-x)^{2}+4\lambda/(1-\lambda).$ (29)
The poles are proportional to $\xi$ and can be rescaled to obtain dimensionless units. For all the considered values of the parameters, $|\mathrm{Re}[\epsilon^{\text{s1}}_{1}]|<|\mathrm{Re}[\epsilon^{\text{s1}}_{2}]|$. The pole labeled as 1 is identified as the QP pole and the one labeled as 2 is the satellite, thus, $Z^{\text{MPA-}G}_{\text{s1}}=Z^{\text{s1}}_{1}$.
@@
Figure 8: Real and imaginary parts of the scaled poles (a), (c) and residues (b), (d) of the Green’s function model of Eq. (27) as a function of the $\lambda$ parameter, where we have fixed its imaginary part to a small value of $\mathrm{Im}~{}\lambda=0.01$. The solid and dashed horizontal gray lines and the black dotted line in (b) are analogous to Fig. 7.
@@
We first focus on the correlation effects by setting $x=0$. Notice that $\lambda=0$ corresponds to $S=0$ (zero correlation), while $\lambda\to 1$ corresponds to the limit of infinite correlation. Figure 7 shows the real and imaginary parts of the scaled poles $\epsilon^{\text{s1}}_{p}/\xi$ [Figs. 7(a) and 7(c)], and the residues $Z_{p}^{\text{s1}}$ [Figs. 7(b) and 7(d)] of $G^{\text{s1}}$ as functions of $\mathrm{Re}~{}\lambda$, for $\mathrm{Im}~{}\lambda=0$ (blue curves) and $\mathrm{Im}~{}\lambda=0.3$ (orange curves). Since $x=0$, $\mathrm{Re}~{}\lambda=0$ and $\mathrm{Im}~{}\lambda=0$ also correspond to the independent-particle limit, in which the QP state has a zero self-energy correction ($\epsilon^{\text{s1}}_{1}=0$) and carries the whole spectral weight ($Z^{\text{MPA-}G}_{\text{s1}}=1$). As a consequence, the satellite $\epsilon^{\text{s1}}_{2}=\xi$ vanishes ($Z^{\text{s1}}_{2}=0$).